Please solve the absolute value of 2X-6 minus 2=12
I get 10 which is correct but I also get -2 which isn't correct. What are the answers?
Thank you!

Answer by unlockmath(1688)

Hello,
Let's set up both equations first.
2X-6 minus 2=12
Add 2 to both sides to get:
2x-6=14
And the other equation will be:
2x-6=-14
Solve for x by adding 6 and dividing by 2, we get:
x=10
x=-4
(note: when we add 6 to -14, we get -8 and then divide by 2 to get
-4.
Make sense?
RJ
www.math-unlock.com